OpenStudy (kidthatbro8):

NEEDS HELP MEDAL & FAN & WHATEVER ELSE What is the product of 5the square root of 3. • 4the square root of 15.? Simplify if possible. 60the square root of 5. 20the square root of 5. 60the square root of 15. 20the square root of 15.

OpenStudy (anonymous):

I'm not 100% sure here, but wouldn't you split the square root of 15 to 5 & 3? So like: \[\sqrt[5]{3}\times \sqrt[4]{15}\] \[\sqrt[5]{3}\times \sqrt[4]{3}\times \sqrt[4]{5}\] the 3s cancel, leaving the 5. and the 5&4 multiply leaving it looking like \[\sqrt[20]{5}\] as the answer. Hope this helps (: